OpenStudy (anonymous):

Suppose that two fair dice are rolled and that the two numbers recorded are multiplied to obtain a final score. Construct and plot the probability mass function and the cumulative distribution function of the final score

OpenStudy (anonymous):

make a six by six multiplication table as a first step

OpenStudy (anonymous):

alright. I got the probabilities of thm.

OpenStudy (anonymous):

then each entry will have probability \(\frac{1}{36}\) add them up to get the probability of each outcome for example, \(P(x=12)=\frac{4}{36}=\frac{1}{9}\) because there are 4 ways to get a 12: \(\{(2,6),(6,2),(3,4),(4,3)\}\)

OpenStudy (anonymous):

So the probability of getting, say, 6 is P(6)=4/36. Thats how you would get the answers for pmf? Then to find cdf, you add up the pdf probabilities?

OpenStudy (anonymous):

ok then that is all a probability mass function is if your random variable is called \(X\) then in this case \(X\) takes on possible values : 1,2,3,4,5,6,8,9,10,12,15,16,18,20,25,30,36 find the probability of each one \[P(X=12)=\frac{1}{9}\] for example the complete answer is the probability mass function

OpenStudy (anonymous):

yes to your second question for example \(F_X(1)=\frac{1}{36}\) and \(F_X(2)=\frac{3}{36}\)

OpenStudy (anonymous):

will the denominator of each P9X) be 36?

OpenStudy (anonymous):

the mass function is sometimes written as \(f_X(x)\) instead of \(P(X=x)\) but they mean the same thing so in your example you have as you said \(f_X(6)=4\)

OpenStudy (anonymous):

oops i meant \[f_X(6)=\frac{4}{36}\] sorry

OpenStudy (anonymous):

ok. and the cdf of 36 has to be 1?

OpenStudy (anonymous):

and yes, the denominators will all be 36 i would not reduce the fraction because for the second part you have to add, so leave the denominators as 36

OpenStudy (anonymous):

the total will be one, yes i.e. \(F_X(36)=1\)

OpenStudy (anonymous):

So after I construct the 6x6 matrix, and the table consisting of X, P(X), and cdf, is that all?

OpenStudy (anonymous):

yes, that is all it asks

OpenStudy (anonymous):

HA! Thank you SO VERY MUCH!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

What do you do for a living? Just a question